With what impulse does a 0.50kg newspaper have to be thrown to give it a velocity of 3 m/s


1
Expert's answer
2021-05-18T11:10:35-0400

"m = 0.5\ kg"

"\vec v_0= 0;"

"\vec v_1= 3\frac{m}{s};"

"\Delta \vec p= m\Delta\vec v= m(\vec v_1-\vec v_0)= 0.5*3 = 1.5\ kg\frac{m}{s}"

"\text{Answer: }1.5\ kg\frac{m}{s}"